OpenStudy (anonymous):

Find the product of the complex numbers. Leave the answer in polar form. Z1=5(cos20º+i sin20º) Z2= 4(cos10º + i sin10º)

Parth (parthkohli):

Can you write these in exponential form?

Parth (parthkohli):

It's a well-known identity:\[\large\cos(\theta) + i\sin (\theta) =e^{i\theta}\]

OpenStudy (anonymous):

i got 20e^(i 30º) = 20(cos 30º + i sin 30º)

terenzreignz (terenzreignz):

Not well known enough, apparently ^_^ To multiply complex numbers in their polar form, then multiply their moduli, and add their arguments... that's the gist of it.

Parth (parthkohli):

@Y$G_Jay Where is 30 degrees?